OpenStudy (yttrium):

Anyone who can give the shortest approach/solution possible will be given a medal. Here's the problem: If a + b = 2^n where n is a real number and f(a) + f(b) = n^2, find f(2002).

OpenStudy (fibonaccichick666):

What class is this for? Theoretical CS?

OpenStudy (lastdaywork):

First, you should try to find some obvious values for f(x); like f(1) = 1/2 f(0) = -1/2 Then, you can break the question using any pattern; I am showing one below - f(2002) + f(46) = 11^2 f(46) + f(18) = 6^2 f(18) + f(14) = 5^2 f(14) + f(2) = 4^2 f(2) + f(0) = 1^2 I think the rest is obvious :)

OpenStudy (anonymous):

How did you get \(f(1)=1/2\)?

OpenStudy (lastdaywork):

Plug-in a = b = 1

OpenStudy (anonymous):

But if \(a = b = 0\), then \(0 + 0 = 2^{-\infty}\)

OpenStudy (lastdaywork):

But n has to be an integer; so we can put a=b=0. So, find f(1) first. Then put a=1 & b=0 :)

OpenStudy (anonymous):

Let \(a=b=2\). Then \(2+2=4=2^2\implies n=2\). \(f(2)+f(2) = 2^2\implies 2f(2) = 4\implies f(2) =2\) now we say \(0+2=2\implies n=1\) \(f(0) + f(2) = 1^2\implies f(0)+2=1\implies f(0)=-1\)

OpenStudy (anonymous):

Consider any number, \(c=2^m\). We can say \(c+c=2^m+2^m = 2^{m+1}\implies n=m+1\) We can say \(f(c)+f(c)=n^2\implies 2f(c) = (m+1)^2\implies f(c) = \frac{(1+m)^2}{2}\)

OpenStudy (lastdaywork):

And that ^^ proves f(x) is not a function in the first place. ;)

OpenStudy (anonymous):

Then \(0+c=2^m\implies n=m\) So \(f(0)+f(c)=n^2\implies f(0) = m^2-\frac{(1+m)^2}{2}\) We can make \(f(0)\) various values based on what \(c\) we use.

OpenStudy (lastdaywork):

I didn't check for inconsistency..my bad XD

OpenStudy (anonymous):

Fibonacci sequence is the same way, unless you add seed data.

OpenStudy (anonymous):

The lack of seed data made me suspicious

OpenStudy (lastdaywork):

What is "seed data" ??

OpenStudy (anonymous):

for example, f(0) = 0 and f(1) =1 for Fibonacci, or f(0) = 1 and f(1) = 1 for factorial

OpenStudy (anonymous):

When you have a recursive function, that is a function defined in terms of itself, then you generally need seed data.

OpenStudy (lastdaywork):

Thanxx :)

OpenStudy (anonymous):

But this question looks so open ended that, I'm not sure if it would work even then.

OpenStudy (anonymous):

For example, is it true that for any number there is pair of \(a,b\) such that \(a+b=2^n\)?

OpenStudy (anonymous):

a unique pair^

OpenStudy (anonymous):

Maybe if you just ignore \(f(0)\) and stop at \(f(2)\) it will work?

OpenStudy (lastdaywork):

The problem is that we'll get different values for f(2) also. For any n; (a,b) is not unique.

OpenStudy (anonymous):

I'm not seeing any obvious connection between f(a), f(b), and a + b = 2^n. Are variables "a" and "b" real numbers?

OpenStudy (anonymous):

Oh, I just had an insight that variable "n" connects them, perhaps?

OpenStudy (dumbcow):

wont f(0) = 0, f(1) = 0 work here as initial values ? 0 + 1 = 2^0 ---> n =0 f(0) + f(1) = 0

ganeshie8 (ganeshie8):

yes, a+b = 2^n => a = 2^n-b f(a) + f(b) = n^2 f(2^n-b) + f(b) = n^2 next, look at LastDayWork's first reply

ganeshie8 (ganeshie8):

^^ that was for cebroski...

OpenStudy (anonymous):

Thanks, @ganeshie8!

ganeshie8 (ganeshie8):

np, also u need to figure out below : 2002 = 2^11 - 46 before looking at LastDayWork's work :)

OpenStudy (lastdaywork):

"LastDayWork's work" XD XD XD

OpenStudy (anonymous):

@dumbcow if \(f(1)=0\) then \(1+1=2^1\implies n=1\) but \(f(1)+f(1)=0+0 =0^2\implies n=0\)

OpenStudy (anonymous):

http://jsfiddle.net/wio_dude/WBpy5/3/ It seems like all initial values are going to be \(c = 2^n\). In these cases, you would use \(f(c) = (1+n)^2/2\). There is no value for \(f(0)\). For cases where \(b \neq 2^n\), then you would find \(n\) such that \(2^{n-1} < b < 2^n\).\[ f(b) = n^2 - f(2^n-b) \]We know that \(2^n-b<b\), because \[\begin{array}{rcl} 2^{n-1}&<&b\\ 2^n &<&2b\\ 2^n-b&<&b \end{array}\]So we know that this recursive function always goes down.

OpenStudy (tkhunny):

If \(a + b = 2^{n}\) And \(f(a) + f(b) = n^{2}\) We have \(f(a) = f(2^{n} - b) = n^{2} - f(b)\) This recursive definition follows exactly LastDaysWork's example sequence.